In the system shown, a horizontal force Fx = 40 N acts on an object of mass m2 = 8.00 kg. A friction force fk = 5 N acts between object m2 and the table top, and m1 = 2.00 kg.
i) Calculate the acceleration of the masses.
ii) Calculate the tension in the string
1
Expert's answer
2020-05-26T12:48:56-0400

Acceleration of the masses


"F_{1} = F_{x}-F_{k}"


"a_{2} = \frac {(40-5)}{8} = 4.375 \frac {m} {sec^{2}}"


"\mu_{2} = \frac {F_{k}} {N} = \frac {5} {8\times9.81}=0.064"


fraction is same "\mu_{2}=\mu{1}=0.064"


Friction force at mass 1 "F_{f} = 0.064\times ((2+8)\times 9.81) =6.28 N"


"a_{1} = \frac {F_{f}} {m_2} = \frac {6.28}{2}=3.14 \frac {m} {sec^2}"


Tension in the string


"T = W_1+W_2 = 2\times 9.81+8\times 9.81 = 98.1 N"